Gulf Shore Lawn and Garden Maintenance provides two general outdoor services: lawn maintenance and garden maintenance. The company charges customers $16.0 per hour for each type of service, but lawn maintenance has higher variable costs ($6.4 per hour) than garden maintenance ($3.2 per hour) because of fuel expenses incurred to operate lawn-mowing equipment. All employees are paid a fixed monthly salary. A contribution format income statement for a recent month for the two services appears below. During the month, 8,000 hours of lawn maintenance services and 2,000 hours of garden maintenance were provided:
Lawn
Maintenance Per
Hour Garden
Maintenance Per
Hour Total
Sales $ 128,000 $ 16.0 $ 32,000 $ 16.0 $ 160,000
Variable expenses 51,200 6.4 6,400 3.2 57,600
Contribution margin $ 76,800 $ 9.6 $ 25,600 $ 12.8 102,400
Fixed expenses 50,176
Operating income $ 52,224
Required:
1. Compute the overall CM ratio for the company.
2. Compute the overall break-even point for the company in sales dollars.
3. Compute the weighted-average CM per hour for the company. (Do not round intermediate calculations. Round your answers to 2 decimal places.)
4. Calculate the overall break-even point for the company in hours.
5. At the overall break-even point in total hours, how many hours of each service must be provided for the company to break-even?
6. Calculate the overall sales in dollars required to earn an after-tax profit of $95,760 if the tax rate is 30%.
Answer #1

Overall CM ratio = Overall contribution margin/overall revenue

= 102400/160,000

=64%

2.overall break even point in sales dollars = fixed costs/contribution margin ratio

= 50176/64%

=$78,400

3.weighted average CM per hour = (16-6.4)*8000/10,000 +(16-3.2)*2000/10,000

=$10.24 per hour

4.break even in hours =fixed costs/contribution margin per hour

= 50176/10.24

= 4900 hours

5.lawn maintenance = 4900*8000/10,000 = 3920 hours

Garden maintenance = 4900*2000/10,000 = 980 hours

6. Dollar sales required =(desired income before tax+fixed cost)/CM ratio

=(95760/0.7 + 50176)/64%

=$292,150